OpenStudy (dls):

Any one please help me with this question or tell me how to go about it ?? Find the interval for "s" so that (3-s)x + sy + (s^2-1)=0 is normal to the curve xy = 4 .

OpenStudy (dls):

@terenzreignz

OpenStudy (dls):

@zepdrix

OpenStudy (amistre64):

hmm, orthogonal functions come to mind

OpenStudy (dls):

\[\LARGE y'=\frac{-(3+s)}{s}\] after differentiating the equation orthogonal functions o.O

OpenStudy (amistre64):

start with derivatives; if the derivatives are perpendicular then they will be orthogonal

OpenStudy (dls):

\[\LARGE y'=\frac{-y}{x}=\frac{x}{y}\]

OpenStudy (dls):

Now i am supposed to equate both but i dont get anything?

OpenStudy (amistre64):

(3-s)x + sy + (s^2-1) = 0 (3-s) + s y' = 0; y' = (s-3)/s ---------------------- xy = 4 y + xy' = 0 y' = -y/x = -4/x^2

OpenStudy (amistre64):

\[\frac{s-3}{s}=-\frac{1}{-4/x^2}\] \[\frac{4}{x^2}(s-3)=s\] \[4(s-3)=sx^2\] etc ...

OpenStudy (dls):

so how do we get the interval??

OpenStudy (amistre64):

notice the since s is in a denominator along the way the s=0 is bad for everyone, thats just an observation

OpenStudy (amistre64):

did yours just goofy too?

OpenStudy (amistre64):

\[4(s-3)=sx^2\] \[4(1-\frac 3s)=x^2\] \[4-\frac {12}s=x^2\] notice that the range of x^2 is from 0 to +inf

OpenStudy (dls):

yes it is

OpenStudy (amistre64):

\[0\le4-\frac{12}s\le\infty\] \[-4\le-\frac{12}s\le\infty\] \[-\infty\le\frac{12}s\le4\] getting any ideas yet?

OpenStudy (amistre64):

\[\frac{12}s\le 4\] \[\frac{12}4\le s\]